# Breaking Single-Tester Limits: Multi-Agent LLMs for Multi-User Feature Testing

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Some research (Kumaresen et al., 2020; Ramírez-Méndez et al., 2021; Ahlgren et al., 2021b, 2020) has introduced the concept of multi-agent systems for software testing, where each agent is an entity designed to autonomously perceive and interact with its environment, and solve problems cooperatively. However, these systems tend to rely on hard-coded rules specific to individual apps, limiting their generalizability across different software. In the last few years, AI multi-agents have shown human behavior patterns (Wang et al., 2023b), and even show superhuman performance in games like AlphaStar for Starcraft (Vinyals et al., 2019), and OpenAI Five for Dota 2 (Berner et al., 2019). The emerging Large Language Models (LLMs) such as GPT (Brown et al., 2020), Claude (web, 2024k), and Mistral (web, 2024l) make it even harder by providing a backbone of autonomous agents that can interact with each other like humans (Park et al., 2023). Inspired by the role-playing capability of LLMs, we propose a novel multi-agent system powered by the LLMs to autonomously conduct multi-user interactive feature testing, with agents assuming distinct, human-like roles for covering multi-user interactive features.
@@
  • We propose a novel approach, MAdroid, that harnesses a multi-agent LLMs framework for the autonomous automation of multi-user interactive tasks.
@@ 2.1. Definition of Multi-User Interactive Feature
Definition: Multi-user interactive feature is a component of software that enables 1multiple users to 2interact with each other in 3nearly real-time within a software.
The definition of multi-user interactive features is dissected into three fundamental elements. 1 The feature enables user-to-user communication through the software. It can be a two-way or multi-way interaction between end-users. Note that interactions between a user and the software or the administrator are not categorized under multi-user interactive features. An example of this would be activity organization, where an end-user books a reservation and the admin confirms it. Such interactions are predominantly administrative and do not embody the peer-to-peer collaborative or communicative intent that is the hallmark of multi-user interactive features. 2 A critical aspect of the interactive feature is its fundamental interactivity. This characteristic is marked by the presence of interactive cues including notifications or pop-up messages, that differ significantly from collaborative features like text editing. These cues play a crucial role in ensuring users are aware of others’ requests and encouraging them to act or respond. 3 The interactive feature pertains to a timely nature. Interactions should occur with minimal delay, though not necessarily instantaneously, to maintain a fluid information exchange and keep users actively engaged. For instance, the call will automatically expire after 59 seconds in Figure 1, which the current automated testing frameworks fail to proceed to further communication activities, potentially impeding testing coverage.
@@ 3.1. Coordinator
The initial stage of our approach is to utilize the Large Language Models (LLMs) to understand, analyze, organize, and strategize for prospective events derived from a high-level interactive task description ($task$). An example of the prompts given to LLMs as Coordinator is detailed in Figure 3. In detail, it first commences by interpreting the instructions as shown in Figure 3-A, which acts as a preliminary framework guiding the LLMs within the task coordination workflow. Subsequently, the agent assumes the role of autonomous planner, which involves figuring out the number of required users, partitioning the task into each user, and determining the task order. Note that our task planning is executed step by step, aligning with the chain-of-thought paradigm (Feng and Chen, 2024).
@@ 3.1.2. Segmenting task
Upon determining the requisite number of users for the task, we then need to segment the $task$ into smaller $sub$-$tasks$ that can be distributed across these interactive users. This involves a potential plan of the task to identify discrete components that can be independently executed by different devices. For example, consider the task of “A sends a watch-together invitation to B, and B accepts the invitation” in Figure 3. It will be broken down into sub-tasks, including “A invites B to watch together” for $user_{A}$ and “B accepts the invitation from A” for $user_{B}$. The detail of the prompt is shown in Figure 3.
@@ 3.2.3. Prompting LLMs
Considering the allocated $sub$-$task$, the range of possible actions, and the current GUI screen, we prompt the LLMs to infer a single viable action to execute on the screen, thereby advancing one step toward the completion of the $task$. An illustrative example of such prompting is provided in Figure 4. Due to the robustness of LLMs, the prompt sentence does not need to adhere strictly to grammar rules (Brown et al., 2020). For instance, given the sub-task to send an invitation to watch together, the output of the action to be performed on the current GUI screen would be [tap] [watch together]. The Operator then executes the suggested actions on the device by using the Android Debug Bridge (ADB) (web, 2024a). This cycle of prompting LLMs for a plausible action is repeated until the LLMs recognize the need to interact with a different user, at which point infer a transition command to another Operator, i.e., [switch] in Figure 4. When an Operator concludes that the $sub$-$task$ has been fully executed without the need for cross-user interaction, it generates a signal indicating task completion (i.e., [end_task]).
@@ 4.1. RQ1: Performance of MAdroid
To gather a collection of multi-user tasks, we recruited four Master students with two-year backgrounds in Android development. These annotators were recruited via the university’s internal Slack channel. To guarantee accurate annotations, we first offered them an introduction to our study along with a sample set of task descriptions that had been annotated by the previous work (Hu et al., 2018; Burns et al., 2022). Subsequently, we required annotators to complete and pass an assessment test before proceeding with the task annotation.
The four annotators were then assigned to investigate the 50 popular apps on Google Play, chosen specifically for their popularity, as it often correlates with a higher likelihood of featuring multi-user interactive tasks. We asked them to independently navigate through these apps, identify multi-user interactive tasks, and label the ground truth, which included task descriptions and the corresponding GUI execution traces. To aid in this process, we also provided them with the apps’ descriptions from Google Play, which could offer additional insights into the apps’ features. Note that we asked the annotators to set up the tasks involved to two or three users for the brevity of the experiments. After the initial labeling, the annotators convened to meet and discuss any differences in their findings, working collaboratively until they reached a consensus on the set of multi-user interactive tasks.
To ensure the quality and practicality of the dataset, we further invited two professional app developers from a large tech company to review the collected tasks. In total, we obtained 41 multi-user interactive tasks derived from 16 apps, with an average of 2.87 actions required per user to complete each task. Note that, all the tasks were freshly identified and labeled by human annotators specifically for this study, mitigating the potential bias for data leakage that could arise from the use of LLMs.
@@
According to the advantages of Coordinator and Observer, our approach MAdroid shows a promising capability for automating multi-user interactive tasks, yet some tasks remain to fail. The primary cause for the failures is related to timing constraints within the interactions. Some multi-user interactive features may require a very quick response (i.e., $<$ 2 seconds). For instance, the Tiktok-2 task of responding to the scrolling comments in Figure 8(a); by the time the LLMs have completed its inference, the comment may have already disappeared from the view. We are optimistic that this challenge may be mitigated by employing more advanced LLMs or by utilizing local LLMs for quicker inference, thereby reducing delays and improving the success rate of task automation. Second, multi-user interactive features can span across multiple apps. For instance, sharing a meeting link via an external app in GoogleMeet-1 in Figure 8(b), requires interaction with not just the meeting app but also the third-party app. Our approach is limited by the extent to which it requires these back-and-forth interactions across different apps.
@@ 4.3. RQ3: Usefulness of MAdroid
Results. Our MAdroid successfully completes 88.3% of the 111 instances of task automation for regression testing. Of the instances that failed, 11 (84.6%) of these are identified as interactive bugs. To further validate the effectiveness of the bugs, we cross-reference them with related issues reported in their issue-tracking repositories on GitHub, where 7 bugs have been confirmed/fixed. For the remaining 4 bugs, which have not been reported earlier, we submit new issue reports under review, and as of now, none have been dismissed. Table 4 provides an overview of the bugs [^1] Due to the anonymity policy, we omit the bug reports we submitted.. Upon further examination of the bugs, we discover that all of them require multi-user interaction to be triggered. Figure 9 illustrates an example of a common multi-user interactive bug where one user invites another to a call, but the user does not receive it due to an app crash, hindering the automated process failure. However, the testing capabilities of multi-user interactive features are not found in the existing automated tools, indicating the usefulness of our approach.
@@ 6.2. Multi-agent Software Testing
Agents have been increasingly applied to support the automation of different software testing activities. One such approach relates to test management (Malz and Jazdi, 2010; Arora and Bhatia, 2018), which aims at selecting an appropriate set of test cases to be executed in every software test cycle using test unit agent and fuzzy logic. Nevertheless, the testing of certain complex software features, such as chat functions, multi-player gaming, and interactions between users with varying levels of access, cannot be effectively conducted by a single-agent. To address this limitation, a body of research (Kumaresen et al., 2020; Ramírez-Méndez et al., 2021; Ahlgren et al., 2020, 2021a) has explored the concept of multi-agent-oriented software testing. This involves the creation of multiple intelligent agents specifically designed to tackle the testing of these intricate features. For instance, Tang (Tang, 2010) introduced a variety of agents, each specialized in different roles equipped with hard-coded capabilities, including test design, execution, and evaluation. Similarly, Dhavachelvan et al. (Dhavachelvan et al., 2006; Dhavachelvan and Uma, 2005) have crafted several ad-hoc agents with predefined patterns to test specific features within the software. However, these approaches rely on hard-coded rules tailored to individual software, which poses challenges for generalization across different software.
@@ 7. Conclusion
In the future, we plan to improve MAdroid in two aspects. First, considering that multi-user interactive features are often time-sensitive, we aim to boost the efficiency of our approach by implementing open-source LLMs on the local server, hence reducing latency and improving response times. Second, since interactive features can extend over various applications, we will explore the potential to facilitate multi-user cross-app interactions.

### v3

@@ Abstract.
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3764535isbn: 979-8-4007-2025-3/26/04ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Some research (Kumaresen et al., 2020; Ram´ırez-Méndez et al., 2021; Ahlgren et al., 2021b, 2020) has introduced the concept of multi-agent systems for software testing, where each agent is an entity designed to autonomously perceive and interact with its environment, and solve problems cooperatively. However, these systems tend to rely on hard-coded rules specific to individual apps, limiting their generalizability across different software. In the last few years, AI multi-agents have shown human behavior patterns (Wang et al., 2023b), and even show superhuman performance in games like AlphaStar for Starcraft (Vinyals et al., 2019), and OpenAI Five for Dota 2 (Berner et al., 2019). The emerging Large Language Models (LLMs) such as GPT (Brown et al., 2020), Claude (web, 2024k), and Mistral (web, 2024l) make it even harder by providing a backbone of autonomous agents that can interact with each other like humans (Park et al., 2023). Inspired by the role-playing capability of LLMs, we propose a novel multi-agent system powered by the LLMs to autonomously conduct multi-user interactive feature testing, with agents assuming distinct, human-like roles for covering multi-user interactive features.
@@
  • We propose a novel approach, MAdroid, that harnesses a multi-agent LLMs framework for the autonomous automation of multi-user interactive tasks. The source code, dataset, and experimental results are in appendix [^1] https://github.com/sidongfeng/MAdroid.
@@ 2.1. Definition of Multi-User Interactive Feature
Definition: Multi-user interactive feature is a component of software that enables multiple users to interact with each other in nearly real-time within a software.
The definition of multi-user interactive features is dissected into three fundamental elements. ①The feature enables user-to-user communication through the software. It can be a two-way or multi-way interaction between end-users. Note that interactions between a user and the software or the administrator are not categorized under multi-user interactive features. An example of this would be activity organization, where an end-user books a reservation and the admin confirms it. Such interactions are predominantly administrative and do not embody the peer-to-peer collaborative or communicative intent that is the hallmark of multi-user interactive features. ②A critical aspect of the interactive feature is its fundamental interactivity. This characteristic is marked by the presence of interactive cues including notifications or pop-up messages, that differ significantly from collaborative features like text editing. These cues play a crucial role in ensuring users are aware of others’ requests and encouraging them to act or respond. ③The interactive feature pertains to a timely nature. Interactions should occur with minimal delay, though not necessarily instantaneously, to maintain a fluid information exchange and keep users actively engaged. For instance, the call will automatically expire after 59 seconds in Figure 1, which the current automated testing frameworks fail to proceed to further communication activities, potentially impeding testing coverage.
@@ 3.1. Coordinator
The initial stage of our approach is to utilize the Large Language Models (LLMs) to understand, analyze, organize, and strategize for prospective events derived from a high-level interactive task description (tasktask). An example of the prompts given to LLMs as Coordinator is detailed in Figure 3. In detail, it first commences by interpreting the instructions as shown in Figure 3-A, which acts as a preliminary framework guiding the LLMs within the task coordination workflow. Subsequently, the agent assumes the role of autonomous planner, which involves figuring out the number of required users, partitioning the task into each user, and determining the task order. Note that our task planning is executed step by step, aligning with the chain-of-thought paradigm (Feng and Chen, 2024).
@@ 3.1.2. Segmenting task
Upon determining the requisite number of users for the task, we then need to segment the tasktask into smaller subsub-taskstasks that can be distributed across these interactive users. This involves a potential plan of the task to identify discrete components that can be independently executed by different devices. For example, consider the task of “A sends a watch-together invitation to B, and B accepts the invitation” in Figure 3. It will be broken down into sub-tasks, including “A invites B to watch together” for $user_{A}$ and “B accepts the invitation from A” for $user_{B}$. The detail of the prompt is shown in Figure 3.
@@ 3.2.3. Prompting LLMs
Considering the allocated subsub-tasktask, the range of possible actions, and the current GUI screen, we prompt the LLMs to infer a single viable action to execute on the screen, thereby advancing one step toward the completion of the tasktask. An illustrative example of such prompting is provided in Figure 4. Due to the robustness of LLMs, the prompt sentence does not need to adhere strictly to grammar rules (Brown et al., 2020). For instance, given the sub-task to send an invitation to watch together, the output of the action to be performed on the current GUI screen would be [tap] [watch together]. The Operator then executes the suggested actions on the device by using the Android Debug Bridge (ADB) (web, 2024a). This cycle of prompting LLMs for a plausible action is repeated until the LLMs recognize the need to interact with a different user, at which point infer a transition command to another Operator, i.e., [switch] in Figure 4. When an Operator concludes that the subsub-tasktask has been fully executed without the need for cross-user interaction, it generates a signal indicating task completion (i.e., [end_task]).
@@ 4.1. RQ1: Performance of MAdroid
To gather a collection of multi-user tasks, four authors with at least two-year backgrounds in Android development were participated. The four annotators were then assigned to investigate the 50 popular apps on Google Play, chosen specifically for their popularity, as it often correlates with a higher likelihood of featuring multi-user interactive tasks. We asked them to independently navigate through these apps, identify multi-user interactive tasks, and label the ground truth, which included task descriptions and the corresponding GUI execution traces. To aid in this process, we also provided them with the apps’ descriptions from Google Play, which could offer additional insights into the apps’ features. Note that we asked the annotators to set up the tasks involved to two or three users for the brevity of the experiments. After the initial labeling, the annotators convened to meet and discuss any differences in their findings, working collaboratively until they reached a consensus on the set of multi-user interactive tasks.
To ensure the quality and practicality of the dataset, we further had two professional app developers from a large tech company to review the collected tasks. In total, we obtained 41 multi-user interactive tasks derived from 16 apps, with an average of 2.87 actions required per user to complete each task. Note that, all the tasks were freshly identified and labeled by human annotators specifically for this study, mitigating the potential bias for data leakage that could arise from the use of LLMs.
@@
According to the advantages of Coordinator and Observer, our approach MAdroid shows a promising capability for automating multi-user interactive tasks, yet some tasks remain to fail. The primary cause for the failures is related to timing constraints within the interactions. Some multi-user interactive features may require a very quick response (i.e., << 2 seconds). For instance, the Tiktok-2 task of responding to the scrolling comments in Figure 8(a); by the time the LLMs have completed its inference, the comment may have already disappeared from the view. We are optimistic that this challenge may be mitigated by employing more advanced LLMs or by utilizing local LLMs for quicker inference, thereby reducing delays and improving the success rate of task automation. Second, multi-user interactive features can span across multiple apps. For instance, sharing a meeting link via an external app in GoogleMeet-1 in Figure 8(b), requires interaction with not just the meeting app but also the third-party app. Our approach is limited by the extent to which it requires these back-and-forth interactions across different apps.
@@ 4.3. RQ3: Usefulness of MAdroid
Results. Our MAdroid successfully completes 88.3% of the 111 instances of task automation for regression testing. Of the instances that failed, 11 (84.6%) of these are identified as interactive bugs. To further validate the effectiveness of the bugs, we cross-reference them with related issues reported in their issue-tracking repositories on GitHub, where 7 bugs have been confirmed/fixed, as shown in Table 4. For the remaining 4 bugs, which have not been reported earlier, we submit new issue reports under review, and as of now, none have been dismissed. Upon further examination of the bugs, we discover that all of them require multi-user interaction to be triggered. Figure 9 illustrates an example of a common multi-user interactive bug where one user invites another to a call, but the user does not receive it due to an app crash, hindering the automated process failure. However, the testing capabilities of multi-user interactive features are not found in the existing automated tools, indicating the usefulness of our approach.
@@ 5. Threats To Validity
Another factor that could affect representativeness is our use of GPT-4 as the LLM model. MAdroid is designed to be model-agnostic, with a primary focus on prompt engineering. Recent studies (Feng et al., 2025; Feng and Chen, 2024) usually adopt a single representative model (e.g., the state-of-the-art GPT-4) to evaluate their model-agnostic methods. Our current experiments with GPT-based models establish a strong initial benchmark and demonstrate the potential of LLM-driven approaches for multi-user interactive feature testing. Evaluating the performance of MAdroid with other LLMs is planned as an extension toward more comprehensive experimentation in future work.
@@ 6.2. Multi-agent Software Testing
Agents have been increasingly applied to support the automation of different software testing activities. One such approach relates to test management (Malz and Jazdi, 2010; Arora and Bhatia, 2018), which aims at selecting an appropriate set of test cases to be executed in every software test cycle using test unit agent and fuzzy logic. Nevertheless, the testing of certain complex software features, such as chat functions, multi-player gaming, and interactions between users with varying levels of access, cannot be effectively conducted by a single-agent. To address this limitation, a body of research (Kumaresen et al., 2020; Ram´ırez-Méndez et al., 2021; Ahlgren et al., 2020, 2021a) has explored the concept of multi-agent-oriented software testing. This involves the creation of multiple intelligent agents specifically designed to tackle the testing of these intricate features. For instance, Tang (Tang, 2010) introduced a variety of agents, each specialized in different roles equipped with hard-coded capabilities, including test design, execution, and evaluation. Similarly, Dhavachelvan et al. (Dhavachelvan et al., 2006; Dhavachelvan and Uma, 2005) have crafted several ad-hoc agents with predefined patterns to test specific features within the software. However, these approaches rely on hard-coded rules tailored to individual software, which poses challenges for generalization across different software.
@@ 7. Conclusion
In the future, we plan to improve MAdroid in three aspects. First, considering that multi-user interactive features are often time-sensitive, we aim to boost the efficiency of our approach by implementing open-source LLMs on the local server, hence reducing latency and improving response times. Second, since interactive features can extend over various applications, we will explore the potential to facilitate multi-user cross-app interactions. Third, although our current work focuses on peer-to-peer and small-group interactions for tractability and experimental control, future research will investigate scaling to more complex social dynamics.
Acknowledgements.
We sincerely appreciate the support of the Bytedance team in facilitating discussions, validating data, and conducting experiments. This research was partially supported by OpenAI Researcher Access Program.